OpenStudy (anonymous):

inverse of y=2x^4, y=4-x^2

myininaya (myininaya):

not 1 to 1 => no inverse

OpenStudy (anonymous):

so what are the restricted domains

myininaya (myininaya):

yes perhaps if you were given some restricted domain we could say we have an inverse

OpenStudy (anonymous):

so what is the restricted domain of the original problem

myininaya (myininaya):

i don't what does it say?

myininaya (myininaya):

and also if you did have a restriction on the domain such that the function is 1 to 1 then the inverse you have above is incorrect

OpenStudy (anonymous):

there is \[2x^{4} and 4-2x^{2} \]

myininaya (myininaya):

2x^4 and 4-2x^2 are not inverses for each other over any restricted domain

myininaya (myininaya):

y=2x^4 is not 1 to 1 does not have inverse

myininaya (myininaya):

she/he never provided us with domain restriction

OpenStudy (anonymous):

how do your restrict the domain of \[27x^{3}\]

myininaya (myininaya):

does it ask you to choose a restriction such that function is 1 to 1?

myininaya (myininaya):

27x^3 is 1 to 1

myininaya (myininaya):

\[y=27x^3 => \frac{y}{27}=x^3 => \sqrt[3]{\frac{y}{27}}=x=> f^{-1}(x)=\sqrt[3]{\frac{x}{27}}\]